SUsing Gauss’ Theorem calculate the flux of the vector field F =
xi ˆ+yj^+z kˆ through
the surface of a cylinder of radius A and height H, which has its axis along the z-axis
and the base of the cylinder is on the xy-plane.
1
Expert's answer
2019-03-14T17:47:02-0400

In this case, the vector flux Ф = divF*V = (1 + 1 + 1)*3.14*A2*H = 9.4*A2*H.